# De novo Design of Polymer Electrolytes with High Conductivity using GPT-based and Diffusion-based Generative Models

### 11 Dec 2023 v1

@@ 1 Introduction
In this study, we delve deep into the intricacies of different advanced generative models for polymer generation. To train the generative models, we leverage a previously curated dataset (HTP-MD dataset) consisting of 6024 different amorphous polymer electrolytes with their ion transport properties computed from MD simulations [10, 31]. In our current investigation, we compare the performance of different generative models’ architectures in (Fig. 1) Unconditional generation, in which the model is trained to learn the “language” of polymers, enabling it to produce novel, chemically valid, unique, and synthesizable polymers. 2) Conditional generation, in which we aim to steer the generative model towards generating candidates exhibiting high ionic conductivity along with other metrics defined for unconditional generation. To ensure high-quality polymer generation, we compare two categories of advanced generative models: one based on a type of large language model (LLM) known as generative pretraining transformer (GPT) and the other based on diffusion models evaluated with multiple metrics. Hyperparameter optimization is undertaken to optimize the performances of these generative models. Through our proposed methodology, we can easily produce a large number of monomer units for polymer electrolytes, which can significantly enrich the current dataset. Furthermore, the top 45 candidates from conditional generation with high ionic conductivity are validated using full-atom MD simulations. Among these candidates, we’ve identified 17 polymer units whose ionic conductivity surpassed all existing data in our trainset, and multiple polymers increased the optimal conductivity value in the trainset by more than 100%. Finally, given the limited data amount and highly constrained polymer space of the HTP-MD dataset, we apply a “pretraining & fine-tuning” strategy to enhance the model’s performance [32]. By employing this strategy, the model attains accelerated convergence and improved data adaptability. In a concurrent study [40], we introduce an advanced polymer discovery framework that leverages the top-performing model from this research. The framework enhances polymer identification through a generative model and a feedback loop with MD simulations, proven by the discovery of novel polymer electrolytes with significantly improved ionic conductivity.
@@ 2.2 Unconditional generation
In addition, we further compare the performances of these models by assessing their ability to replicate distributions of various polymer properties, including the length of the SMILES string, conductivity, and transference number. Among these properties, the transference number is the fraction of the total ionic charge carried by a particular ion species in the electrolyte. In this work, we refer to the transference number specifically as the cation transference number. As for polymer electrolytes, a higher ionic conductivity and cation transference number will lead to more efficient and safer battery operation[47]. Given these properties, we examined how well the distribution of the generated set aligns with the training set (Fig. 2d). To calculate the ionic conductivity and transference number, a previously developed graph neural network (GNN)[31] which is trained on the HTP-MD dataset is implemented. As shown in the figures, the minGPT model (blue lines in figures) reproduces the distribution accurately for all three properties. In contrast, the other two diffusion-based models (green & red lines) tend to produce narrower distributions for various properties compared to the training data. The quantitative evaluation is based on the difference calculation of two distributions using Kullback-Leibler (KL) divergence.
@@ 2.3 Conditional generation
We utilize the optimal model architectures obtained from hyperparameter tuning of unconditional generation tasks to produce polymer electrolytes with high conductivities. The GNN model is employed to predict the ionic conductivity of generated polymers. Fig. 3c displays the conductivity distribution, comparing the generated set with the training set, which reveals a distinct shift toward the high-conductivity domain for all generated sets from the three models. Once again, we plot the model comparison using the six metrics (Fig. 3d). Aligning with the findings from the unconditional generation, the minGPT model surpasses both the 1Ddiffusion and diffusion-LM models, achieving a superior average score. The top five candidates from the generated sets of the three models are presented in Fig. 3b. In contrast to the diverse structures of unconditionally generated monomer units, the conditionally generated polymers with the highest conductivities exhibit consistent patterns (Fig. 3b). For example, majority of the generated polymers possess linear backbones and contain the “-O-CH2-CH2-“ fragment within their chains. Given the high ionic conductivity of PEO-like polymers, such an observation is to be expected.
@@ 2.4 Molecular dynamics validation
With MD simulations as a validation tool, a closed-loop framework can be established. Each iteration within this framework comprises a sequence of steps: conditional generation of polymers, preliminary screening via GNN predictions, evaluation through MD simulations, and data augmentation incorporating new findings. In a separate study [40], we have explored this opportunity by presenting a robust polymer discovery framework based on improving generative models’ performance through iteratively incorporating MD simulation feedback. We have shown this approach can facilitate the generation of promising polymer electrolyte candidates originating from the refinement of trained machine learning models. We envision integrating this approach with high-throughput synthesis methodologies in the future to pave the way for the development of a self-driving laboratory.
@@ 3 Discussion
In this study, we compare and benchmark advanced deep generative AI models to learn the complex language of polymers, enabling the generation of promising electrolyte candidates. Our proposed methodology significantly enriches the relatively undeveloped data pool of polymer electrolytes. We test two mainstream generative models including GPT and diffusion-based models. The GPT-based models outperform diffusion-based architectures with better performance and lower training/inference costs. The results of hyperparameter tuning indicate that small models can also show great performances given the limited number of data we have. We leverage multiple evaluation metrics proposed previously to provide a comprehensive evaluation of the generation task. Given the positive and negative correlations between these metrics, we further examine the model performance in a multi-objective setup. Moreover, by conditioning the generative process, our models are capable of producing polymers tailored to specific target properties. Empowered by our proposed approach, the generation of novel, unique, chemically valid, and potentially synthesizable polymers can be accomplished in microseconds. This significantly surpasses the capabilities of human chemists, offering a dramatic acceleration in the materials discovery process. By applying this method, we have already discovered multiple polymer candidates that surpass all existing data from our HTP-MD dataset.
@@
By integrating MD simulations as a validation tool, we have established a comprehensive closed-loop workflow in a parallel study [40]. This workflow will encompass a series of sequential steps: generative design of new polymers, calculating properties of generated polymer candidates via MD simulations which serves as a validation technique, and the enrichment of our database with these new findings. This iterative loop is capable of continuously generating new electrolyte candidates and progressively refining the developed models. Furthermore, the scope of this workflow extends beyond electrolyte material discovery. Considering the vast array of applications for polymer materials, this method shows tremendous potential in revolutionizing various polymer systems. It could significantly impact sectors such as biomedical devices, sustainable packaging solutions, advanced aerospace components, and high-performance textiles, among others. Each of these fields could benefit from the enhanced efficiency and targeted material properties this workflow promises.
@@
Another contribution of this study is showing how LLMs and other large deep generative models can be utilized for the generative design of polymers under conditions of limited labeled data availability. The pivotal process in this context is pretraining, which effectively maximizes the potential of these models and enables them to learn from broader contexts before being fine-tuned for specific applications. This pretraining and fine-tuning strategy empowers large foundational models to surpass traditional generative models like GANs or VAEs, in domain-specific and design-focused tasks. Moreover, the ability of these models even with sparse datasets is crucial, especially when applying this workflow to experimental data, which is often scarce, unstructured, and expensive to generate.
@@ 4.3 Hyperparameter selection
Apart from the training step, other hyperparameters are used for different models and both selected and optimal values of these hyperparameters are listed in Table S3. The other two hyperparameters of the minGPT model are model architecture and temperature. The "model architecture" corresponds to the HuggingFace’s Transformer-based model architecture [60]. Meanwhile, “temperature,” a characteristic feature of GPT-style models usually ranging from 0 to 1, dictates the degree of randomness incorporated into the generation process. Notably, a higher temperature leads to more random, diverse, and creative text generation, but it also increases the likelihood of producing incoherent or irrelevant outputs, often termed “hallucinations.” As for the 1Ddiffusion model, dim is the initial hidden dimension of 1D UNet and dim_mult determines the scaling of layer dimensions. For instance, with dim_mult set to (1, 2), the hidden dimension of the second layer is twice that of the first layer. Finally, when it comes to diffusion-LM models, model architecture is again the HuggingFace’s model. The in_channel is the input channel dimensions of embeddings.

### 07 Mar 2024 v2

@@ 1 Introduction
In this study, we delve deep into the intricacies of different advanced generative models for polymer generation. To train the generative models, we leverage a previously curated dataset (HTP-MD dataset) consisting of 6024 different amorphous polymer electrolytes with their ion transport properties computed from MD simulations [32, 10]. In our current investigation, we compare the performance of different generative models’ architectures in (Fig. 1): 1) Unconditional generation, in which the model is trained to learn the “language” of polymers, enabling it to produce novel, chemically valid, unique, and synthesizable polymers. 2) Conditional generation, in which we aim to steer the generative model towards generating candidates exhibiting high ionic conductivity along with other metrics defined for unconditional generation. To ensure high-quality polymer generation, we compare two categories of advanced generative models: one based on a minimal implementation of a large language model (LLM) known as generative pretraining transformer (GPT) and the other based on diffusion models evaluated with multiple metrics. Hyperparameter optimization is undertaken to optimize the performances of these generative models. Through our proposed methodology, we can easily produce a large number of monomer units for polymer electrolytes, which can significantly enrich the current dataset. Furthermore, the top 45 candidates from conditional generation with high ionic conductivity are validated using full-atom MD simulations. Among these candidates, we’ve identified 17 polymer units whose ionic conductivity surpassed all existing data in our trainset, and multiple polymers increased the optimal conductivity value in the train set by more than 100%. Finally, given the limited data amount and highly constrained polymer space of the HTP-MD dataset, we apply a “pretraining & fine-tuning” strategy to enhance the model’s performance [33]. By employing this strategy, the model attains accelerated convergence and improved data adaptability. In a concurrent study [41], we introduce an advanced polymer discovery framework that leverages the top-performing model from this research. The framework enhances polymer identification through an active feedback loop which augments the dataset on the fly with generated high-performance polymers validated by MD simulations. This self-improving framework further allows the fast discovery of novel polymer electrolytes with significantly improved ionic conductivity.
@@ 2.2 Unconditional generation
In addition, we further compare the performances of these models by assessing their ability to replicate distributions of various polymer properties, including the length of the SMILES string, conductivity, and transference number. Among these properties, the transference number is the fraction of the total ionic charge carried by a particular ion species in the electrolyte. In this work, we refer to the transference number specifically as the cation transference number. As for polymer electrolytes, a higher ionic conductivity and cation transference number will lead to more efficient and safer battery operation[48]. Given these properties, we examined how well the distribution of the generated set aligns with the training set (Fig. 2d). To calculate the ionic conductivity and transference number, we utilized a graph neural network (GNN) which was developed in our previous study[32]. The GNN model was also trained on the HTP-MD dataset and outperformed classical ML algorithms like random forest. As shown in the figures, the minGPT model (blue lines in figures) reproduces the distribution accurately for all three properties. In contrast, the other two diffusion-based models (green & red lines) tend to produce narrower distributions for various properties compared to the training data. The quantitative evaluation is based on the difference calculation of two distributions using Kullback-Leibler (KL) divergence.
@@ 2.3 Conditional generation
We utilize the optimal model architectures obtained from hyperparameter tuning of unconditional generation tasks to produce polymer electrolytes with high conductivities. The GNN model proposed in our previous study[32] is employed to predict the ionic conductivity of generated polymers. Fig. 3c displays the conductivity distribution, comparing the generated set with the training set, which reveals a distinct shift toward the high-conductivity domain for all generated sets from the three models. Once again, we plot the model comparison using the six metrics (Fig. 3d). Aligning with the findings from the unconditional generation, the minGPT model surpasses both the 1Ddiffusion and diffusion-LM models, achieving a superior average score. The top five candidates from the generated sets of the three models are presented in Fig. 3b. In contrast to the diverse structures of unconditionally generated monomer units, the conditionally generated polymers with the highest conductivities exhibit consistent patterns (Fig. 3b). For example, majority of the generated polymers possess linear backbones and contain the “-O-CH2-CH2-“ fragment within their chains. Given the high ionic conductivity of PEO-like polymers, such an observation is to be expected.
@@ 2.4 Molecular dynamics validation
With MD simulations as a validation tool, a closed-loop framework can be established. Each iteration within this framework comprises a sequence of steps: conditional generation of polymers, preliminary screening via our ML predictions, evaluation through MD simulations, and data augmentation incorporating new findings. In a separate study [41], we have explored this opportunity by presenting a robust polymer discovery framework based on improving generative models’ performance through iteratively incorporating MD simulation feedback. We have shown this approach can facilitate the generation of promising polymer electrolyte candidates originating from the refinement of trained machine learning models. We envision integrating this approach with high-throughput synthesis methodologies in the future to pave the way for the development of a self-driving laboratory.
@@ 3 Discussion and Conclusion
In this study, we compare and benchmark advanced deep generative AI models to learn the complex language of polymers, enabling the generation of promising electrolyte candidates. Our proposed methodology significantly enriches the relatively undeveloped data pool of polymer electrolytes. We test two mainstream generative models including GPT and diffusion-based models. The GPT-based models outperform diffusion-based architectures with better performance and lower training/inference costs. The results of hyperparameter tuning indicate that small models can also show great performances given the limited number of data we have. We design a systematic evaluation scheme for polymer generation with multiple metrics, and leverage it to comprehesively evaluate the generation tasks. Given the positive and negative correlations between these metrics, we further examine the model performance in a multi-objective setup. Moreover, by conditioning the generative process, our models are capable of producing polymers tailored to specific target properties. Empowered by our proposed approach, the generation of novel, unique, chemically valid, and potentially synthesizable polymers can be accomplished in microseconds. This significantly surpasses the capabilities of human chemists, offering a dramatic acceleration in the materials discovery process. By applying this method, we have already discovered multiple polymer candidates that surpass all existing data from our HTP-MD dataset.
@@
By integrating MD simulations as a validation tool, we have established a targeted polymer discovery framework with comprehensive closed-loop workflow in a parallel study [41]. This workflow will encompass a series of sequential steps: generative design of new polymers, calculating properties of generated polymer candidates via MD simulations which serves as a validation technique, and the enrichment of our database with these new findings. This iterative loop is capable of continuously generating new electrolyte candidates and progressively refining the developed models. Furthermore, the scope of this workflow extends beyond electrolyte material discovery. Considering the vast array of applications for polymer materials, this method shows tremendous potential in revolutionizing various polymer systems. It could significantly impact sectors such as biomedical devices, sustainable packaging solutions, advanced aerospace components, and high-performance textiles, among others. Each of these fields could benefit from the enhanced efficiency and targeted material properties this workflow promises.
@@
Another contribution of this study is showing how LLMs and other large deep generative models can be utilized for the generative design of polymers under conditions of limited labeled data availability. The pivotal process in this context is pretraining, which effectively maximizes the potential of these models and enables them to learn from broader contexts before being fine-tuned for specific applications. Although the capacity of LLMs is not fully exploited when we train the model with our small HTP-MD dataset, the implementation of LLMs can really benefit the generation when pretrained on a vast array of diverse polymers. This pretraining and fine-tuning strategy empowers large foundational models to surpass traditional generative models like GANs or VAEs, in domain-specific and design-focused tasks. Moreover, the ability of these models, even with sparse datasets, is crucial, especially when applying this workflow to experimental data, which is often scarce, unstructured, and expensive to generate.
@@ 4.3 Hyperparameter selection
Apart from the training step, other hyperparameters are used for different models and both selected and optimal values of these hyperparameters are listed in Table S3. The other two hyperparameters of the minGPT model are model architecture and temperature. The “model architecture” corresponds to the HuggingFace’s Transformer-based model architecture [61]. Meanwhile, “temperature,” a characteristic feature of GPT-style models usually ranging from 0 to 1, dictates the degree of randomness incorporated into the generation process. Notably, a higher temperature leads to more random, diverse, and creative text generation, but it also increases the likelihood of producing incoherent or irrelevant outputs, often termed “hallucinations.” As for the 1Ddiffusion model, dim is the initial hidden dimension of 1D UNet and dim_mult determines the scaling of layer dimensions. For instance, with dim_mult set to (1, 2), the hidden dimension of the second layer is twice that of the first layer. Finally, when it comes to diffusion-LM models, model architecture is again the HuggingFace’s model. The in_channel is the input channel dimensions of embeddings. The detailed results for the hyperparameter grid search for three models are listed in Tables S4, S5, and S6. To summarize the findings presented in the tables, the impact of enhancing the model’s complexity, either through deeper architecture or more intricate hyperparameter configurations, parallels the effect of increasing training steps. Such modifications typically enhance the model’s validity and its ability to synthesize data. However, they may also lead to a reduction in novelty and uniqueness.